OpenStudy (anonymous):

The length of the longer leg of a right triangle is 16 more than four times the length of the shorter leg. The length of the hypotenuse is 17 more than four times the length of the shorter leg. Find the side lengths of the triangle

zepdrix (zepdrix):



zepdrix (zepdrix):

16 MORE THAN 4 times, another way we can write this is 4times the thing, plus 16

zepdrix (zepdrix):

And then uhhh, similar for the hypotenuse, yah?

OpenStudy (anonymous):

well yes the hypotenuse would be 4x+17 so my equation would be x^2+(4x+16)^2=(4x+17)^2 or would it be x+(4x+16)+(4x+17)=180

zepdrix (zepdrix):

Applying the Pythagorean Theorem,\[\Large\rm x^2+(4x+16)^2=(4x+17)^2\]Yah the first one looks good :) Since it's a formula we apply to `side lengths`. The other one you listed, we apply to `angles`, which we don't have here.

OpenStudy (anonymous):

well good to know that i eventually got it right but i think i did it wrong on my test :( oh well better i learn it now than not know at all thank you very much!

zepdrix (zepdrix):

c: